# Restricted Spectral Gap Decomposition for Simulated Tempering Targeting Mixture Distributions

### v1

@@ Appendix A Simulated Tempering Metropolis–Hastings algorithm
Input: function $f$ and inverse temperature sequence $\beta_{1}<\cdots<\beta_{L}$
@@
     Repeat Algorithm 1 $s$ times to get samples $(x_{j})_{j=1}^{s}$.
@@ B.1. Proof of Lemma 1
which is symmetric with respect to $i$ and $i^{\prime}$. ∎
@@ B.2. Proof of Theorem 2
The $[L]\times\mathcal{X}^{0}$-restricted Dirichlet form $\mathcal{E}_{[L]\times\mathcal{X}^{0}}$ of the simulated tempering Markov chain $M$, defined in Definition 2, can be expressed by
@@
(B.3) ¯$\overline{\mathcal{E}}(\overline{g},\overline{g})=\overline{\mathcal{E}}^{J}(\overline{g},\overline{g})+\overline{\mathcal{E}}^{I}(\overline{g},\overline{g}),$
@@
¯$\overline{\mathcal{E}}^{J}(\overline{g},\overline{g})$ $=\frac{1}{2\theta}\sum_{i=1}^{L}\sum_{j,j^{\prime}=1}^{n}(\overline{g}(i,j)-\overline{g}(i,j^{\prime}))^{2}\,r_{i}w_{(i,j)}P_{(i,j)}(\mathcal{X}^{0})\,\overline{M}((i,j),(i,j^{\prime})),$
¯$\overline{\mathcal{E}}^{I}(\overline{g},\overline{g})$ $=\frac{1}{2\theta}\sum_{j=1}^{n}\sum_{i,i^{\prime}\in[L]\colon i^{\prime}=i\pm 1}(\overline{g}(i,j)-\overline{g}(i^{\prime},j))^{2}\,r_{i}w_{(i,j)}P_{(i,j)}(\mathcal{X}^{0})\,\overline{M}((i,j),(i^{\prime},j)).$
@@
(B.4) ¯$\overline{\mathcal{E}}^{J}(\overline{g},\overline{g})\leq\frac{2(1-\lambda)}{\theta}\sum_{i=1}^{L}\sum_{j=1}^{n}\frac{r_{i}w_{(i,j)}}{P_{(i,j)}(\mathcal{X}^{0})}\;\mathrm{Var}_{P_{(i,j)},\mathcal{X}^{0}}(g_{i}).$
@@
(B.5) ¯$\overline{\mathcal{E}}^{J}(\overline{g},\overline{g})$ $=\frac{1-\lambda}{2\theta}\sum_{i=1}^{L}\sum_{j,j^{\prime}=1}^{n}(\overline{g}(i,j)-\overline{g}(i,j^{\prime}))^{2}r_{i}w_{(i,j)}\int_{\mathcal{X}^{0}}p_{(i,j)}(x)p_{i}(j^{\prime}|x)\,\mathrm{d}x$
@@
(B.12) ¯$\overline{\mathcal{E}}^{I}(\overline{g},\overline{g})\leq\frac{3\lambda}{2\theta}\sum_{i=1}^{L}\sum_{j=1}^{n}\frac{r_{i}w_{(i,j)}}{P_{(i,j)}(\mathcal{X}^{0})}\;\mathrm{Var}_{P_{(i,j)},\mathcal{X}^{0}}(g_{i})+\frac{3\lambda}{2\theta}\mathcal{E}^{I}_{\mathcal{X}^{0}}(g,g).$
@@
(B.13) ¯$\overline{\mathcal{E}}^{I}(\overline{g},\overline{g})$ $=\frac{\lambda}{4\theta}\sum_{i,i^{\prime}\in[L]\colon i^{\prime}=i\pm 1}\sum_{j=1}^{n}(\overline{g}(i,j)-\overline{g}(i^{\prime},j))^{2}r_{i}w_{(i,j)}\int_{\mathcal{X}^{0}}p_{(i,j)}(x)a((i,j,x),(i^{\prime},j,x))\mathrm{d}x$
(B.14) $\leq\frac{3\lambda}{4}{\mathsf{E}}_{\tilde{P}}\Bigl{[}\bigl{(}\overline{g}(I,J)-g(I,X)\bigr{)}^{2}+\bigl{(}g(I,X)-g(I^{\prime},X)\bigr{)}^{2}+\bigl{(}\overline{g}(I^{\prime},J)-g(I^{\prime},X)\bigr{)}^{2}\Bigr{]},$
@@
Recall that ¯$\overline{\mathcal{E}}(\overline{g},\overline{g})=\overline{\mathcal{E}}^{J}(\overline{g},\overline{g})+\overline{\mathcal{E}}^{I}(\overline{g},\overline{g}).$ By Lemma 7,
(B.19) ¯$\overline{\mathcal{E}}^{J}(\overline{g},\overline{g})\leq\frac{2(1-\lambda)}{\theta}\sum_{i=1}^{L}\sum_{j=1}^{n}\frac{r_{i}w_{(i,j)}}{P_{(i,j)}(\mathcal{X}^{0})}\;\mathrm{Var}_{P_{(i,j)},\mathcal{X}^{0}}(g_{i})\leq\frac{2(1-\lambda)C_{1}C_{2}}{\theta\phi}\sum_{i=1}^{L}r_{i}\mathcal{E}_{i,\mathcal{X}^{0}}(g_{i},g_{i}).$
@@
¯$\overline{\mathcal{E}}^{I}(\overline{g},\overline{g})$ $\leq\frac{3\lambda}{2\theta}\sum_{i=1}^{L}\sum_{j=1}^{n}\frac{r_{i}w_{(i,j)}}{P_{(i,j)}(\mathcal{X}^{0})}\;\mathrm{Var}_{P_{(i,j)},\mathcal{X}^{0}}(g_{i})+\frac{3\lambda}{2\theta}\mathcal{E}^{I}_{\mathcal{X}^{0}}(g,g)$
$\leq\frac{3\lambda C_{1}C_{2}}{2\theta\phi}\sum_{i=1}^{L}r_{i}\mathcal{E}_{i,\mathcal{X}^{0}}(g_{i},g_{i})+\frac{3\lambda}{2\theta}\mathcal{E}^{I}_{\mathcal{X}^{0}}(g,g)$
@@
$\frac{1}{\theta^{2}}\operatorname{Var}_{P,[L]\times\mathcal{X}^{0}}(g)=\operatorname{Var}_{P_{0}}(g)\leq\frac{3\lambda C_{3}}{2\theta}\mathcal{E}^{I}_{\mathcal{X}^{0}}(g,g)+\frac{C_{1}C_{2}\left[(4-\lambda)C_{3}+2\right]}{2\theta\phi}\sum_{i=1}^{L}r_{i}\mathcal{E}_{i,\mathcal{X}^{0}}(g_{i},g_{i}).$
@@
$\operatorname{Var}_{P,[L]\times\mathcal{X}^{0}}(g)$ $\leq\max\left\{\frac{3}{2}\theta C_{3},\;\frac{\theta C_{1}C_{2}}{2\phi(1-\lambda)}\left({(4-\lambda)C_{3}}+2\right)\right\}{\mathcal{E}_{[L]\times\mathcal{X}^{0}}}(g,g)$
@@ B.3. Proof for the Mixing Times
We first recall the mixing time bound given in Atchadé et al. (2011) using restricted spectral gaps.
Lemma 9 (Atchadé et al. (2011)).
@@ C.1. Comparison of STMH chain with approximate STMH chain
To compare the STMH chain defined in Definition 4 with the approximate STMH chain in Definition 5, it suffices to compare the stationary density $p^{*}_{i}$ with $\widetilde{p}_{i}$ and transition kernel $M^{*}_{i}$ with $\widetilde{M}_{i}$. For the former, we use Lemma 7.3 of Ge et al. (2018), which shows that varying the temperature is roughly the same as changing the variance of a Gaussian distribution. For the latter, we derive a bound in Lemma 11.
Lemma 10 (Lemma 7.3 of Ge et al. (2018)).
@@
From now on, we assume that $\mathcal{X}^{0}\subseteq\mathbb{R}^{d}$ is a measurable subset. Our next result, Lemma 12, shows that it suffices to obtain a lower bound on the $[L]\times\mathcal{X}^{0}$-restricted spectral gap of the approximate STMH chain in order to derive a corresponding bound for the STMH chain.
@@
Let $M^{*}$ be the STMH chain defined in Definition 4, and let $\widetilde{M}$ be the approximate STMH chain defined in Definition 5. Assume that the mixture weights satisfy $w_{\min}:=\min_{1\leq j\leq n}w_{j}>0$. Then, the $[L]\times\mathcal{X}^{0}$-restricted spectral gaps of $M^{*}$ and $\widetilde{M}$ satisfy the inequality
@@
for every $(i,x)$, and the same inequality holds for $p^{*}(i,x)$ and $\widetilde{p}(i,x)$ since the weights $(r_{i})_{i=1}^{L}$ are the same for $P^{*}$ and $\widetilde{P}$. Let $\mathcal{E}^{*}_{[L]\times\mathcal{X}^{0}},\widetilde{\mathcal{E}}_{[L]\times\mathcal{X}^{0}}$ denote the ${[L]\times\mathcal{X}^{0}}$-restricted Dirichlet forms associated with $M^{*}$ and $\widetilde{M}$ respectively. Fix a function $g\in\mathcal{L}^{2}([L]\times\mathcal{X},\widetilde{p})$ and define $g_{i}(x):=g(i,x)$ for each $(i,x)$. By Definition 1, it suffices to show that
@@
(C.5) ~$\widetilde{\mathcal{E}}_{[L]\times\mathcal{X}^{0}}(g,g)=(1-\lambda)\sum_{i=1}^{L}r_{i}\,\widetilde{\mathcal{E}}_{i,\mathcal{X}^{0}}(g_{i},g_{i})+\lambda\,\widetilde{\mathcal{E}}^{I}_{\mathcal{X}^{0}}(g,g),$
@@
~$\widetilde{\mathcal{E}}_{i,\mathcal{X}^{0}}(g_{i},g_{i})$ $=\frac{1}{2}\,\int_{\mathcal{X}^{0}}\!\int_{\mathcal{X}^{0}}\bigl{(}g_{i}(x)-g_{i}(y)\bigr{)}^{2}\,\widetilde{p}_{i}(x)\,\widetilde{M}_{i}(x,\mathrm{d}y)\,\mathrm{d}x$
@@
~$\widetilde{\mathcal{E}}^{I}_{\mathcal{X}^{0}}(g,g)$ $=\frac{1}{4}\sum_{i,i^{\prime}\in[L]\,:\,i^{\prime}=i\pm 1}\int_{\mathcal{X}^{0}}\bigl{(}g_{i}(x)-g_{i^{\prime}}(x)\bigr{)}^{2}\min\left\{r_{i}\widetilde{p}_{i}(x),\,r_{i^{\prime}}\widetilde{p}_{i^{\prime}}(x)\right\}\,\mathrm{d}x$
@@ C.2. Restricted Spectral Gap of the Approximate STMH Chain
where $R>0$ is a fixed radius. To obtain a lower bound on the $[L]\times\mathcal{X}^{0}$-restricted spectral gap of the approximate STMH chain, we invoke Theorem 2. Assumption 1 holds by our construction of $\tilde{P}$. The following lemmas verify the other two assumptions required for this corollary.
@@
where ~$\widetilde{\mathcal{E}}_{(i,j),\mathcal{X}^{0}}$ denotes the $\mathcal{X}^{0}$-restricted Dirichlet form of the kernel $\widetilde{M}_{(i,j)}$, and ~$\widetilde{\mathcal{E}}_{i,\mathcal{X}^{0}}$ denotes the $\mathcal{X}^{0}$-restricted Dirichlet form of the kernel $\widetilde{M}_{i}$, as defined in Definition 5.
@@
For any nonnegative real sequences $\{a_{j}\}$ and $\{b_{j}\}$, we have the inequality $\min\left\{\sum_{j}a_{j},\;\sum_{j}b_{j}\right\}\;\geq\;\sum_{j}\min\left\{a_{j},\,b_{j}\right\}.$ Applying this to $\widetilde{p}_{i}=\sum_{j=1}^{n}w_{j}\widetilde{p}_{(i,j)}$, we obtain
@@
~$\widetilde{\mathcal{E}}_{i,\mathcal{X}^{0}}(g_{i},g_{i})$ $=\frac{1}{2}\int_{\mathcal{X}^{0}}\int_{\mathcal{X}^{0}}\left(g_{i}(x)-g_{i}(z)\right)^{2}q(x,z)\min\left\{\widetilde{p}_{i}(x),\widetilde{p}_{i}(z)\right\}\,\mathrm{d}x\,\mathrm{d}z$
@@
We lower bound the $\mathcal{X}^{0}$-restricted spectral gap of each Metropolis–Hastings chain $\widetilde{M}_{(i,j)}$ using the path method of Yuen (2000) in the following lemma.
@@
We use the linear path method described in Section 2 of Yuen (2000). This approach also extends to the restricted spectral gap setting; see, for example, Atchadé (2021) and Chang and Zhou (2024) where the canonical path method has been adapted to the restricted spectral gap in discrete spaces. For each pair $(x,y)\in\mathcal{X}^{0}\times\mathcal{X}^{0}$, we construct a linear path connecting $x$ to $y$, with all intermediate points lying in $\mathcal{X}^{0}$. Fix a step size $\delta>0$, and define the number of steps along the path by
@@
where $\widetilde{m}_{(i,j)}$ denotes the transition density corresponding to the kernel $\widetilde{M}_{(i,j)}$, and $q(u,v)$ is the Gaussian proposal density associated with kernel $\widetilde{M}_{(i,j)}$. As shown in Section 2 of Yuen (2000), the set of paths $\Gamma$ satisfies the regularity conditions and, for any $(u,v)\in\gamma_{xy}$, the associated Jacobian satisfies $J_{x,y}(u,v)=b_{xy}^{d}$ (see (Yuen, 2000, page 5) for details). Then, by Theorem 2.1 and Corollary 2.4 in Yuen (2000), we have
@@
$A=\operatorname*{ess\,sup}_{(u,v)\in E}\left\{\frac{1}{T_{(i,j)}(u,v)}\sum_{\gamma_{xy}\ni(u,v)}|\gamma_{xy}|\,\widetilde{p}_{(i,j)}(x)\,\widetilde{p}_{(i,j)}(y)\,b_{xy}^{d}\right\}.$
where $|\gamma_{xy}|$ denotes the length of the path $\gamma_{xy}$. Since $\tilde{p}_{(i,j)}$ is log-concave, for any $(u,v)\in\gamma_{xy}$, we have

### v2

@@ 5. Concluding Remarks
6. Acknowledgments
Jhanvi Garg and Quan Zhou were supported in part by NSF grants DMS-2245591 and DMS-2311307. Krishnakumar Balasubramanian was supported in part by NSF grant DMS-2413426.
@@ Appendix A Simulated Tempering Metropolis–Hastings algorithm
Input: function $f$, inverse temperature sequence $\beta_{1}<\cdots<\beta_{L}$ and number of samples $s$.
@@
  Repeat Algorithm 1 until $s$ samples $(x_{j})_{j=1}^{s}$ are obtained at temperature level $\ell$.
@@
Remark 7.
Algorithm 1 is always run for a fixed number of steps $N$ and returns the sample obtained at the final step. In Algorithm 2, if this sample is not from the desired temperature level, Algorithm 1 is simply re-run for another NN steps.
@@ B.1. Proof of Lemma 1
which is symmetric with respect to $i$ and $i^{\prime}$.
@@ B.2. Proof of Theorem 2
The $L]\times\mathcal{X}^{0}-$restricted Dirichlet form $\mathcal{E}_{[L]\times\mathcal{X}^{0}}$ of the simulated tempering Markov chain $M$, defined in Definition 2, can be expressed by
@@
(B.3) $\overline{\mathcal{E}}(\overline{g},\overline{g})=\overline{\mathcal{E}}^{J}(\overline{g},\overline{g})+\overline{\mathcal{E}}^{I}(\overline{g},\overline{g}),$
@@
$\overline{\mathcal{E}}^{J}(\overline{g},\overline{g})$ $=\frac{1}{2\theta}\sum_{i=1}^{L}\sum_{j,j^{\prime}=1}^{n}(\overline{g}(i,j)-\overline{g}(i,j^{\prime}))^{2}\,r_{i}w_{(i,j)}P_{(i,j)}(\mathcal{X}^{0})\,\overline{M}((i,j),(i,j^{\prime})),$
$\overline{\mathcal{E}}^{I}(\overline{g},\overline{g})$ $=\frac{1}{2\theta}\sum_{j=1}^{n}\sum_{i,i^{\prime}\in[L]\colon i^{\prime}=i\pm 1}(\overline{g}(i,j)-\overline{g}(i^{\prime},j))^{2}\,r_{i}w_{(i,j)}P_{(i,j)}(\mathcal{X}^{0})\,\overline{M}((i,j),(i^{\prime},j)).$
@@
(B.4) $\overline{\mathcal{E}}^{J}(\overline{g},\overline{g})\leq\frac{2(1-\lambda)}{\theta}\sum_{i=1}^{L}\sum_{j=1}^{n}\frac{r_{i}w_{(i,j)}}{P_{(i,j)}(\mathcal{X}^{0})}\;\mathrm{Var}_{P_{(i,j)},\mathcal{X}^{0}}(g_{i}).$
@@
(B.5) $\overline{\mathcal{E}}^{J}(\overline{g},\overline{g})$ $=\frac{1-\lambda}{2\theta}\sum_{i=1}^{L}\sum_{j,j^{\prime}=1}^{n}(\overline{g}(i,j)-\overline{g}(i,j^{\prime}))^{2}r_{i}w_{(i,j)}\int_{\mathcal{X}^{0}}p_{(i,j)}(x)p_{i}(j^{\prime}|x)\,\mathrm{d}x$
@@
(B.12) $\overline{\mathcal{E}}^{I}(\overline{g},\overline{g}){\leq\frac{3\lambda}{\theta}\sum_{i=1}^{L}\sum_{j=1}^{n}\frac{r_{i}w_{(i,j)}}{P_{(i,j)}(\mathcal{X}^{0})}\;\mathrm{Var}_{P_{(i,j)},\mathcal{X}^{0}}(g_{i})+\frac{3\lambda}{\theta}\mathcal{E}^{I}_{\mathcal{X}^{0}}(g,g)}.$
@@
(B.13) $\overline{\mathcal{E}}^{I}(\overline{g},\overline{g})$ $=\frac{\lambda}{4\theta}\sum_{i,i^{\prime}\in[L]\colon i^{\prime}=i\pm 1}\sum_{j=1}^{n}(\overline{g}(i,j)-\overline{g}(i^{\prime},j))^{2}r_{i}w_{(i,j)}\int_{\mathcal{X}^{0}}p_{(i,j)}(x)a((i,j,x),(i^{\prime},j,x))\mathrm{d}x$
(B.14) ${\leq\frac{3\lambda}{2}{\mathsf{E}}_{\tilde{P}}\Bigl[\bigl(\overline{g}(I,J)-g(I,X)\bigr)^{2}+\bigl(g(I,X)-g(I^{\prime},X)\bigr)^{2}+\bigl(\overline{g}(I^{\prime},J)-g(I^{\prime},X)\bigr)^{2}\Bigr]},$
@@
Recall that $\overline{\mathcal{E}}(\overline{g},\overline{g})=\overline{\mathcal{E}}^{J}(\overline{g},\overline{g})+\overline{\mathcal{E}}^{I}(\overline{g},\overline{g}).$ By Lemma 7,
(B.19) $\overline{\mathcal{E}}^{J}(\overline{g},\overline{g})\leq\frac{2(1-\lambda)}{\theta}\sum_{i=1}^{L}\sum_{j=1}^{n}\frac{r_{i}w_{(i,j)}}{P_{(i,j)}(\mathcal{X}^{0})}\;\mathrm{Var}_{P_{(i,j)},\mathcal{X}^{0}}(g_{i})\leq\frac{2(1-\lambda)C_{1}C_{2}}{\theta\phi}\sum_{i=1}^{L}r_{i}\mathcal{E}_{i,\mathcal{X}^{0}}(g_{i},g_{i}).$
@@
$\overline{\mathcal{E}}^{I}(\overline{g},\overline{g})$ ${\leq\frac{3\lambda}{\theta}\sum_{i=1}^{L}\sum_{j=1}^{n}\frac{r_{i}w_{(i,j)}}{P_{(i,j)}(\mathcal{X}^{0})}\;\mathrm{Var}_{P_{(i,j)},\mathcal{X}^{0}}(g_{i})+\frac{3\lambda}{\theta}\mathcal{E}^{I}_{\mathcal{X}^{0}}(g,g)}$
${\leq\frac{3\lambda C_{1}C_{2}}{\theta\phi}\sum_{i=1}^{L}r_{i}\mathcal{E}_{i,\mathcal{X}^{0}}(g_{i},g_{i})+\frac{3\lambda}{\theta}\mathcal{E}^{I}_{\mathcal{X}^{0}}(g,g)}$
@@
${\frac{1}{\theta^{2}}\operatorname{Var}_{P,[L]\times\mathcal{X}^{0}}(g)=\operatorname{Var}_{P_{0}}(g)\leq\frac{3\lambda C_{3}}{\theta}\mathcal{E}^{I}_{\mathcal{X}^{0}}(g,g)+\frac{C_{1}C_{2}\left[(2+\lambda)C_{3}+1\right]}{\theta\phi}\sum_{i=1}^{L}r_{i}\mathcal{E}_{i,\mathcal{X}^{0}}(g_{i},g_{i})}.$
@@
$\operatorname{Var}_{P,[L]\times\mathcal{X}^{0}}(g)$ ${\leq\max\left\{3\theta C_{3},\;\frac{\theta C_{1}C_{2}}{\phi(1-\lambda)}\left({(2+\lambda)C_{3}}+1\right)\right\}{\mathcal{E}_{[L]\times\mathcal{X}^{0}}}(g,g)}$
@@ B.3. Proof for the Mixing Times
We first recall the mixing time bound given in Atchadé et al. [2011] using restricted spectral gaps.
Lemma 9 (Atchadé et al. [2011]).
@@ C.1. Comparison of STMH chain with approximate STMH chain
To compare the STMH chain defined in Definition 4 with the approximate STMH chain in Definition 5, it suffices to compare the stationary density $p^{*}_{i}$ with $\widetilde{p}_{i}$ and transition kernel $M^{*}_{i}$ with $\widetilde{M}_{i}$. For the former, we use Lemma 7.3 of Ge et al. [2018], which shows that varying the temperature is roughly the same as changing the variance of a Gaussian distribution. For the latter, we derive a bound in Lemma 11.
Lemma 10 (Lemma 7.3 of Ge et al. [2018]).
@@
From now on, we assume that $\mathcal{X}^{0}\subseteq\mathbb{R}^{d}$ is a measurable subset. Our next result, Lemma 12, shows that it suffices to obtain a lower bound on the $L]\times\mathcal{X}^{0}-$restricted spectral gap of the approximate STMH chain in order to derive a corresponding bound for the STMH chain.
@@
Let $M^{*}$ be the STMH chain defined in Definition 4, and let $\widetilde{M}$ be the approximate STMH chain defined in Definition 5. Assume that the mixture weights satisfy $w_{\min}:=\min_{1\leq j\leq n}w_{j}>0$. Then, the $L]\times\mathcal{X}^{0}-$restricted spectral gaps of $M^{*}$ and $\widetilde{M}$ satisfy the inequality
@@
[…]
@@
(C.5) $\widetilde{\mathcal{E}}_{[L]\times\mathcal{X}^{0}}(g,g)=(1-\lambda)\sum_{i=1}^{L}r_{i}\,\widetilde{\mathcal{E}}_{i,\mathcal{X}^{0}}(g_{i},g_{i})+\lambda\,\widetilde{\mathcal{E}}^{I}_{\mathcal{X}^{0}}(g,g),$
@@
$\widetilde{\mathcal{E}}_{i,\mathcal{X}^{0}}(g_{i},g_{i})$ $=\frac{1}{2}\,\int_{\mathcal{X}^{0}}\!\int_{\mathcal{X}^{0}}\bigl(g_{i}(x)-g_{i}(y)\bigr)^{2}\,\widetilde{p}_{i}(x)\,\widetilde{M}_{i}(x,\mathrm{d}y)\,\mathrm{d}x$
@@
$\widetilde{\mathcal{E}}^{I}_{\mathcal{X}^{0}}(g,g)$ $=\frac{1}{4}\sum_{i,i^{\prime}\in[L]\,:\,i^{\prime}=i\pm 1}\int_{\mathcal{X}^{0}}\bigl(g_{i}(x)-g_{i^{\prime}}(x)\bigr)^{2}\min\left\{r_{i}\widetilde{p}_{i}(x),\,r_{i^{\prime}}\widetilde{p}_{i^{\prime}}(x)\right\}\,\mathrm{d}x$
@@ C.2. Restricted Spectral Gap of the Approximate STMH Chain
where $R>0$ is a fixed radius. To obtain a lower bound on the $L]\times\mathcal{X}^{0}-$restricted spectral gap of the approximate STMH chain, we invoke Theorem 2. Assumption 1 holds by our construction of $\widetilde{P}$. The following lemmas verify the other two assumptions required for this theorem.
@@
where $\widetilde{\mathcal{E}}_{(i,j),\mathcal{X}^{0}}$ denotes the $\mathcal{X}^{0}$-restricted Dirichlet form of the kernel $\widetilde{M}_{(i,j)}$, and $\widetilde{\mathcal{E}}_{i,\mathcal{X}^{0}}$ denotes the $\mathcal{X}^{0}$-restricted Dirichlet form of the kernel $\widetilde{M}_{i}$, as defined in Definition 5.
@@
For any nonnegative real sequences $\{a_{j}\}$ and $\{b_{j}\}$, we have the inequality
$\min\left\{\sum_{j}a_{j},\;\sum_{j}b_{j}\right\}\;\geq\;\sum_{j}\min\left\{a_{j},\,b_{j}\right\}.$
Applying this to $\widetilde{p}_{i}=\sum_{j=1}^{n}w_{j}\widetilde{p}_{(i,j)}$, we obtain
@@
$\widetilde{\mathcal{E}}_{i,\mathcal{X}^{0}}(g_{i},g_{i})$ $=\frac{1}{2}\int_{\mathcal{X}^{0}}\int_{\mathcal{X}^{0}}\left(g_{i}(x)-g_{i}(z)\right)^{2}q(x,z)\min\left\{\widetilde{p}_{i}(x),\widetilde{p}_{i}(z)\right\}\,\mathrm{d}x\,\mathrm{d}z$
@@
We lower bound the $\mathcal{X}^{0}$-restricted spectral gap of each Metropolis–Hastings chain $\widetilde{M}_{(i,j)}$ using the path method of Yuen [2000] in the following lemma.
@@
We use the linear path method described in Section 2 of Yuen [2000]. This approach also extends to the restricted spectral gap setting; see, for example, Atchadé [2021] and Chang and Zhou [2024] where the canonical path method has been adapted to the restricted spectral gap in discrete spaces. For each pair $(x,y)\in\mathcal{X}^{0}\times\mathcal{X}^{0}$, we construct a linear path connecting $x$ to $y$, with all intermediate points lying in $\mathcal{X}^{0}$. Fix a step size $\delta>0$, and define the number of steps along the path by
@@
where $\widetilde{m}_{(i,j)}$ denotes the transition density corresponding to the kernel $\widetilde{M}_{(i,j)}$, and $q(u,v)$ is the Gaussian proposal density associated with kernel $\widetilde{M}_{(i,j)}$. As shown in Section 2 of Yuen [2000], the set of paths $\Gamma$ satisfies the regularity conditions and, for any $(u,v)\in\gamma_{xy}$, the associated Jacobian satisfies $J_{x,y}(u,v)=b_{xy}^{d}$ (see [Yuen, 2000, page 5] for details). Then, by Theorem 2.1 and Corollary 2.4 in Yuen [2000], we have
@@
$A=\operatorname*{ess\,sup}_{(u,v)\in E}\left\{\frac{1}{T_{(i,j)}(u,v)}\sum_{\gamma_{xy}\ni(u,v)}|\gamma_{xy}|\,\widetilde{p}_{(i,j)}(x)\,\widetilde{p}_{(i,j)}(y)\,b_{xy}^{d}\right\},$
and $|\gamma_{xy}|$ denotes the length of the path $\gamma_{xy}$. Since $\tilde{p}_{(i,j)}$ is log-concave, for any $(u,v)\in\gamma_{xy}$, we have